Problem 2-24 Schedule of Cost of Goods Manufactured; Income Statement; Cost Behaviour (LO1, LO2, LO3, LO4, LO5) Carlton ManufRequired: 1-a. Prepare a schedule of cost of goods manufactured of the companys income statement for the year. CARLTON MANUF1-b. Prepare a schedule cost of goods sold section of the companys income statement for the year. Carlton Manufacturing Sche3. Assume that in the following year the company expects to produce 22,600 units. What average cost per unit and total cost wPlease check my answers.
Answer #1

The schedule of cost of goods manufactured and cost of goods sold looks fine.

If 17600 units are produced, then average cost per direct materials = 181280/17600 = 10.3

average cost per rent factory building = 72560/17600 = 4.12

Since material cost is variable, it will change as per the number of units produced. However, rent, factory cost will remain the same i.e 72560.

So average cost per direct material for 22600 units is 10 and average cost per rent, factory building is 72560/22600 = 3.21.

All your answers are fine.

All the best.